An economics club has 30 members.
a) If a committee of 4 is to be selected, in how many ways can the selection be made?
b) In how many ways can a committee of at least 1 and at most 3 be selected?

Answer by stanbon(57214)

An economics club has 30 members.
a) If a committee of 4 is to be selected, in how many ways can the selection be made?
---
30C4 = (30*29*28*17)/(1*2*3*4) = 27405
=============================================

b) In how many ways can a committee of at least 1 and at most 3 be selected?
1 member: 30 ways
2 members: (30*29)/(1*2) = 435
3 members: (30*29*28)/(1*2*3) = 4060
----------------------------------------
Cheers,
Stan H.